Grade 12th passMechanics

If radius of a circular plate is increasing at a rate of 2 cm/s. Find the rate of change of its area when circumference would be 2 cm.

abhishek
9 Years agoGrade 12th pass

1 Answer

ayush
9 Years ago
dr/dt=2cm/s.  when c=2cm, r=1/pi. a=pi*r*r. da/dt=pi* dr^2/dt= pi*2r*dr/dt=pi*1/pi*2*dr/dt=4cm^2/s
ans=4cm^2/sec